What mass of ammonium thiocyanate must be used if it is to react completely with 6.5g barium hydroxide octahydrate?


1
Expert's answer
2021-05-04T14:15:46-0400

Ammonium thiocyanate = NH4SCN

Barium hydroxide octahydrate = Ba(OH)2 · 8H2O


Solution:

Balanced chemical equation:

2NH4SCN + Ba(OH)2 · 8H2O → Ba(SCN)2 + 2NH3 + 10H2O

According to the equation above: n[NH4SCN]/2 = n[Ba(OH)2 · 8H2O]


The molar mass of Ba(OH)2 · 8H2O is 315.46 g/mol.

Hence,

n[Ba(OH)2 · 8H2O] = (6.5 g) × (1 mol / 315.46 g) = 0.020605 mol


n[NH4SCN] = 2 × n[Ba(OH)2 · 8H2O] = 2 × 0.020605 mol = 0.04121 mol

The molar mass of NH4SCN is 76.122 g/mol.

Hence, the mass of NH4SCN will be:

m(NH4SCN) = (0.04121 mol NH4SCN) × (76.122 g NH4SCN / 1 mol NH4SCN) = 3.137 g NH4SCN

Mass of NH4SCN = 3.137 g = 3.14 g


Answer: 3.14 grams of ammonium thiocyanate (NH4SCN) must be used.